Fler Manufacturing has 6 million shares of common stock outstanding. The current share price is $78, and the book value per share is $5. Fller Manufacturing also has two bond issues outstanding. The first bond Issue has a face value of $65 million, has a 7 percent coupon, and sells for 98 percent of par. The second Issue has a face value of $35 million, has a 8 percent coupon, and sells for 106 percent of par. The first Issue matures in 22 years, the second in 5 yeers The most recent dividend was $5 and the dividend growth rate is 8 percent. Assume that the overall cost of debt is the weighted average of that Implied by the two outstanding debt Issues. Both bonds make semiannual payments. The tax rate Is 30 percent. Required: What is the companys WACC? (Do not round your intermediate calculations.) 13.14% o 18% o 6.57% О 13.75% 0 6.88%
Answer #1

Bond 1:

Coupon = 0.07 * 65,000,000 = 4,550,000 / 2 = 2,275,000

Price = 0.98 * 65,000,000 = 63,700,000

Number of periods = 22 * 2 = 44

Yield to maturity = 7.1822%

Keys to use in a financial calculator: 2nd I/Y 2, FV 65,000,000, PV -63,700,000, PMT 2,275,000, N 44, CPT I/Y

Bond 2:

Coupon = 0.08 * 35,000,000 = 2,800,000 / 2 = 1,400,000

Price = 1.06 * 35,000,000 = 37,100,000

Number of periods = 5 * 2 = 10

Yield to maturity = 6.5726%

Keys to use in a financial calculator: 2nd I/Y 2, FV 35,000,000, PV -37,100,000, PMT 1,400,000, N 10, CPT I/Y

Total market value of bonds = 63,700,000 + 37,100,000 = 100,800,000

Weight of bond 1 = 63,700,000 / 100,800,000 = 0.6319

Weight of bond 2 = 37,100,000 / 100,800,000 = 0.3681

Weighted average cost of debt = 0.6319*0.071822 + 0.3681*0.065726

Weighted average cost of debt = 0.045384 + 0.024194

Weighted average cost of debt = 0.069578 or 6.9578%

Cost of equity = ( D1 / share price) + growth rate

Cost of equity = [(5 * 1.08) / 78] + 0.08

Cost of equity = 0.1492 or 14.92%

Market value of equity = 6,000,000 * 78 = 468,000,000

Total market value = 468,000,000 + 100,800,000 = 568,800,000

Weight of equity = 468,000,000 / 568,800,000 = 0.8228

Weight of debt = 100,800,000 / 568,800,000 = 0.1772

WACC = Weight of debt*after tax cost of debt + weight of equity*cost of equity

WACC = 0.1772*0.069578*(1 - 0.3) + 0.8228*0.1492

WACC = 0.00863 + 0.12276

WACC = 0.1314 or 13.14%